Point North Inc. issued bonds on September 1, 2020, with a par value of $180,000. The bonds mature in 15 years and pay 9.50% annual interest in two semiannual payments. On the issue date, the annual market rate of interest for the bonds turned out to be 8%

a)What is the semiannual interest payment for these bonds? (

b)How many semiannual interest payments will be made on these bonds over their life?

c)Calculate the issue price of the bonds on September 1, 2020

d)Present the journal entry that would be made to record the bonds’ issuance
Answer #1
a. Par(maturity value) Semiannual Semi annual cash
rate interest payment
180,000 4.75% = 8550
b. Number of payments = 15*2 = 30

4)

Table values are based on
n= 30
i = 4%
Cash flow Table value Amount Present value
Par (maturity) value 180000 * 0.3083 = 55494
Interest (annuity) 8550 * 17.2920 = 147847
Price of bonds 203341

Journal entry :

Date account and explanation debit credit
Cash 203341
Premium on bonds payable 23341
Bonds payable 180,000
(To record bond issue)